OpenStudy (deadshot):

The length of a day in Philadelphia, Pennsylvania, is given by the sinusoidal function, f(t), which represents the number of hours of daylight t days into the year: f(t)=11+2.2sin((2(pi)/365)t-1.25). How many hours of daylight are there on the twelfth day of the year? Round the answer to the nearest tenth.

OpenStudy (callisto):

Please do not post the same question multiple times in other subgroups. Thank you.

OpenStudy (deadshot):

This one is different from the other one. This one has \[f(t)=11+2.2\sin((2\Pi/365)t-1.25)\] while the other one has \[f(t)=11+2.2\sin((2\Pi/365)t-1.35)\]